OpenStudy (mellisa):

Solve the equation for y. Then find the value of y for each value of x. 2x-5y=19; x=1,0,1 Solve equation for y.

OpenStudy (rhystic):

So you have to bring 5y to the opposite side to make it positive and then 19 to the other side to isolate 5y and then divide both sides by 5 to get y alone. Did you get that much? After that its simply plugging in the values for x.

OpenStudy (rhystic):

alternatively you can subtract 2x from both sides and then divide both sides by -5 which will yield the same results for y.

OpenStudy (mellisa):

im really bad with this stuff.... show me how to solve it?

OpenStudy (rhystic):

so you have 2x - 5y = 19 step 1: 2x - 5y - 2x = 19 - 2x (subtract 2x from both sides) yielding: -5y = 19 - 2x step 2: -5y / -5 = (19-2x) / -5 (divide both sides by -5 to get y alone) yielding: y = (19-2x)/(-5) so now you have y = -(19-2x)/5 . Now all you have to to do is plug in the three values for x to get the three values for y. make sense?

OpenStudy (mellisa):

great thanks Rhystic

OpenStudy (rhystic):

sure thing.